A population of values has a normal distribution with μ=81μ=81 and σ=39σ=39. You intend to draw a random sample of size n=176n=176. State answers to one decimal places.

Find P92, which is the score separating the bottom 92% scores from the top 8% scores.
P92 (for single values) =

Find P92, which is the mean separating the bottom 92% means from the top 8% means.
P92 (for sample means) =
Answer #1

Solution :

Given that ,

mean = \mu =  81

standard deviation = \sigma =  39

a ) P( Z < z) = 92%
   P(Z < z) = 0.92

z = 1.41

Using z-score formula,

x = z * \sigma + \mu

x = 1.41 * 39 + 81

= 135.99

P92 = 136.0

b ) n = 176

\mu\bar x = 81

\sigma\bar x = \sigma / \sqrt n = 39 / \sqrt 176 =  2.9397

P( Z < z) = 92%
   P(Z < z) = 0.92

z = 1.41

Using z-score formula,

\bar x= z * \sigma \bar x+ \mu \bar x

\bar x = 1.41 * 2.9397+ 81

= 85.14

P92 = 85.1